User Tools

Site Tools


lbaops:lbamar2007:v194bholog

xraid disks W8-14.

Disk W8 was not included in the array as it had a problem.

There was no problem during v215b using the same disks.

Hopefully v215b data is recoverable.

Last 10 minutes of this exeriment were wind stowed.

The log from experiment v194b has been copied to: ftp.atnf.csiro.au:/pub/people/vlbi/incoming/

ANTAB and UVFLG files have been copied to the same place.

Observers: Simon/Brett

2007.081.16:04:21.27;“DAS profile mp16s_n.pro

2007.081.16:04:26.98;”Coherence from a tone injected at 842.2 MHz

2007.081.16:04:40.77;“Data being recorded to /data/xraid0_0 (xraids W8-14)

2007.081.16:04:45.71;”Agilent set to 14 GHz, SMY set to 777 MHz

2007.081.16:04:50.02;“DAS straight through to VSIC, bandwidth per channel 16 MHz, recording chans 1,3

2007.081.16:17:07.76;”Disk W8 in pack W8-14 is not operating. We seem to have enough space for the experiment though

2007.081.16:21:28.87;“Last Tsys effected by signal for coherence check

2007.081.23:51:34.59;”coherent

2007.081.23:54:30.25;“weather: warm, cloudy and calm

2007.082.01:58:50.99;”missing 5% of PPS signals.

2007.082.02:26:04.89;“disk W8 has been unhappy for this run.

2007.082.02:26:31.72;”weather: Wind is increasing.

2007.082.02:26:55.11;“weather: cloudy and warm

2007.082.03:34:01.65;”Wind is increasing

2007.082.03:51:05.11;“wind stow. That might be all from Ho for this experiment

Regards,

Brett

lbaops/lbamar2007/v194bholog.txt · Last modified: 2015/12/18 16:38 (external edit)